OpenStudy (anonymous):

PLEASE HELP!!!! WILL GIVE MEDAL AND BE FAN!!!! Parker earned $26 by selling 20 glasses of fresh-squeezed lemonade at his lemonade stand on a Saturday morning. He charged $1.00 for a small glass and $1.50 for a large glass. Use a model to determine how many large glasses he sold. (A) 8 (B) 10 (C) 12 (D) 14

OpenStudy (anonymous):

its C) 12 12 x 1.50 = 18 8 x 1 = 8 18 + 8 = 26

OpenStudy (anonymous):

yeah you can plug in the answers to see what fits, but how do you do this without plugging them in? you create two equations, correct?

OpenStudy (anonymous):

x=small, y=large 26$=1.00x + 1.50y and 20= x+y

OpenStudy (anonymous):

so C is the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnx @Christhewiz187 and @MissLovelyPearlOyster

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

welcome:) do you know how to do the problem, or juss lookin for the answer? it's kay, we all get lazy ;)

OpenStudy (anonymous):

Looking for the answer lol @MissLovelyPearlOyster

OpenStudy (anonymous):

ohh haha okay well thats fine, we all do it. message me when you got another problem :)

OpenStudy (anonymous):

or me lol